OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y -2x + 8? (0, -4) (5, -1) (7, 0) (9, 1)

jimthompson5910 (jim_thompson5910):

What's between y and -2x?

OpenStudy (anonymous):

greater than or equal to sign

jimthompson5910 (jim_thompson5910):

checking (0,-4) y >= -2x + 8 -4 >= -2(0) + 8 ... Plug in x = 0 and y = -4 -4 >= 0 + 8 -4 >= 8 since that's FALSE, this means that (0, -4) is NOT a solution

jimthompson5910 (jim_thompson5910):

Do the same with the remaining points.

OpenStudy (anonymous):

okay thank you!

jimthompson5910 (jim_thompson5910):

np